# Assessing High-Risk Systems: An EU AI Act Verification Framework

### v1

@@ 4.10. R10: Technical Documentation
Article 11 and Annex IV of the AI Act require providers of high-risk AI systems to prepare and maintain comprehensive technical documentation demonstrating compliance with the regulation. This documentation must provide sufficient detail to assess the system’s conformity, including its intended purpose, design specifications, data sources, testing methods, and risk management processes. Proper documentation ensures transparency, enables traceability of design choices, and supports both internal governance and external conformity assessment.
Requirement Guidelines Methods Type Target
R10.1: System overview & intended purpose prEN 18286; ISO 9001; ISO/IEC 42001 M10.1 System overview and intended purpose documented in alignment with Annex IV and QMS requirements. C FP
R10.2: Design specifications & architecture prEN 18286; ISO 9001; ISO/IEC 42001 M10.2 Architecture and design specifications captured with complete traceability to risks and requirements. C M, P
R10.3: Dataset description & provenance prEN 18286; ISO 9001; ISO/IEC 42001 M10.3 Dataset provenance, quality attributes and limitations documented in compliance with Art. 10 and Annex IV. C D
R10.4: Performance metrics & validation prEN 18286; ISO 9001; ISO/IEC 42001 M10.4 Performance, robustness and validation evidence compiled and traceable to requirements, risks and intended purpose. T FP
R10.5: Compliance evidence mapping prEN 18286; ISO 9001; ISO/IEC 42001 M10.5 Structured compliance mapping demonstrating how Annex IV and Article 17 obligations are satisfied. C P
@@ 4.11. R11: Record-keeping
Article 18 of the AI Act mandates that providers retain automatically generated logs and relevant records to ensure traceability and accountability throughout the AI lifecycle. Record-keeping supports post-market monitoring, facilitates incident investigation, and provides evidence for conformity assessments. It must include documentation of model versions, dataset lineage, and data retention or deletion policies, consistent with organisational governance and data protection requirements.
Requirement Guidelines Methods Type Target
@@
R11.1: Operational logs prEN 18286; ISO 9001; ISO/IEC 42001 M11.1 Operational logging system in place with records supporting Art. 12 transparency and PMM requirements. C FP
@@
R11.2: Version control records prEN 18286; ISO 9001; ISO/IEC 42001 M11.2 Version control ensures full traceability and reproducibility of AI system states over time. C M
@@
R11.3: Model & dataset lineage prEN 18286; ISO 9001; ISO/IEC 42001 M11.3 Full lineage enables auditing, incident analysis, risk traceability and evidence-based updates. C D, M
@@
R11.4: Retention & deletion policies prEN 18286; ISO 9001; ISO/IEC 42001 M11.4 Retention/deletion policies implemented to ensure compliance, minimise risk and maintain only necessary records. C P
5. Example application
This section illustrates how the proposed framework can be applied in practice through a concrete example drawn from a real, ongoing high-risk AI use case currently being addressed in collaboration with a Swedish automotive manufacturer. While the framework itself is generic and technology-agnostic, this example demonstrates how it can be used to structure assurance activities across the full lifecycle of an AI system, from design and testing to validation and documentation.
The use case concerns an AI-based system for detecting cyberattacks within connected vehicles. Modern vehicles rely on internal communication networks that enable electronic components to exchange information in real time. Protecting these networks against malicious interference is critical, as cyberattacks may compromise both vehicle safety and reliability. To mitigate these risks, automotive manufacturers increasingly deploy Intrusion Detection Systems that monitor in-vehicle communications and identify abnormal or potentially harmful behaviour.
In this setting, the IDS employs AI techniques to analyse network traffic within the vehicle and detect patterns associated with known or emerging attacks. The system processes data from in-vehicle networks, extracts relevant features, and uses machine-learning models to distinguish normal operation from suspicious activity. To support transparency and engineering oversight, the system also integrates explainability mechanisms that allow engineers to understand why specific activities are flagged as intrusions.
Designed for real-time deployment in vehicle control units or central gateways, this real-world use case illustrates how the proposed framework can be applied in a complex, safety-critical automotive context. In particular, the framework enables a structured and traceable assessment of data handling, model behaviour, explainability, and documentation, providing evidence to support safety, cybersecurity, and regulatory requirements.
Legal Requirement (AI Act) Test / Control Type Method Instantiation in the Use Case
@@ 5. Example application
6. Discussion and Conclusion
The framework presented in Section 3 operationalises compliance by establishing explicit correspondences between legal requirements, recognised standards, and assessment protocols. Its stratified architecture—distinguishing controls and testing procedures across data, model, process, and product dimensions—articulates how different modes of assurance interrelate throughout the AI system lifecycle. This structure facilitates systematic self-assessment by delineating which technical and organisational elements require documented evidence of conformity, thereby supporting the structuring of documentation and monitoring activities. Such guidance may prove particularly valuable for organisations lacking established compliance infrastructure.
Within regulatory sandboxes, the framework serves as a common reference architecture for aligning testing protocols and reporting practices with AI Act requirements. By translating legal principles into empirically verifiable criteria, it enables more transparent and structured dialogue between developers and regulatory authorities. Furthermore, it promotes comparability across sandbox experiments and supports coherent assessment of compliance readiness at different stages of system development. For formal conformity assessment, the mapping provides auditors and notified bodies with an instrument for distinguishing procedural from empirical evidence, enhancing traceability, and complementing sector-specific standards through a unified vocabulary for evaluating regulatory conformity.
Beyond immediate applications, the framework contributes to the longer-term objective of partially or fully automated compliance verification. By formalising relationships between legal requirements, control mechanisms, and testing procedures, it establishes a foundation for computational systems capable of selecting, executing, and interpreting appropriate assessment protocols. This aligns with emerging initiatives that translate regulatory obligations into executable testing workflows and interactive monitoring dashboards (buscemi2025sandbox; ki_prueftoolsuite_2025). A natural extension involves developing an ontology that links each legal provision to its associated controls and testable mechanisms, thereby supporting partial automation and interoperability across assessment tools and regulatory databases. Incorporating a temporal dimension would additionally enable longitudinal analysis of how regulatory frameworks, assessment methodologies, and compliance expectations evolve, facilitating both retrospective evaluation and prospective governance planning.
The framework possesses inherent limitations. Any mapping from legal provisions to technical methods necessarily involves interpretive judgement, and the appropriateness of individual assessment procedures remains context-dependent. The deliberately generic level of abstraction may constrain direct application without domain-specific adaptation. Consequently, the framework should be understood not as a prescriptive checklist but as a reference architecture subject to iterative refinement as standards mature and regulatory practice develops. Future research directions include developing prioritisation schemes among requirements, integrating continuous monitoring mechanisms, and extending the framework to encompass additional EU instruments relevant to AI governance.
This study provides an initial architectural foundation for decomposing AI Act obligations into structured, empirically testable components. By establishing coherent mappings among requirements, assessment methods, and implementation guidelines, it seeks to bridge epistemic and operational divides among regulators, risk managers, developers, technical auditors, and prospective certification bodies. The framework is conceived as an evolving infrastructure—supporting transparent, consistent, and traceable compliance verification while fostering evidence-based dialogue between technical and legal communities. It represents both a methodological contribution and an invitation to collaborative development toward a shared, adaptive, and continuously improving ecosystem for trustworthy AI assurance.

### v2

@@ 3.10. R10: Technical Documentation
Article 11 and Annex IV of the AI Act require providers of high-risk AI systems to prepare and maintain comprehensive technical documentation demonstrating compliance with the regulation. This documentation must provide sufficient detail to assess the system’s conformity, including its intended purpose, design specifications, data sources, testing methods, and risk management processes. Proper documentation ensures transparency, enables traceability of design choices, and supports both internal governance and external conformity verification.
Requirement References Methods Type Target
R10.1: System overview & intended purpose ISO/IEC 42001 (§§ 4.3, 6.1.2, 7.5); ISO/IEC 23894 (§  6.3); ISO/IEC 22989 M10.1 System overview and intended purpose documented in alignment with Annex IV and QMS requirements. C FP
R10.2: Design specifications & architecture ISO/IEC 42001(§§  8.1–8.3, 7.5); ISO/IEC 23894 (§§  6.4–6.5) M10.2 Architecture and design specifications captured with complete traceability to risks and requirements. C D&M
R10.3: Dataset description & provenance ISO/IEC 42001 (§§  7.5, 8.3); ISO/IEC 5259-1; ISO 8000-8; ISO/IEC TR 24027 M10.3 Dataset provenance, quality attributes and limitations documented in compliance with Art. 10–11 and Annex IV. C D&M
R10.4: Performance metrics & validation ISO/IEC 42001 (§§  8.3, 9.1, 7.5); ISO/IEC 23894 (§§   6.4–6.5); ISO/IEC 24029-1/-2; ISO/IEC 24970; ISO/IEC 27001; ISO/IEC 27005 M10.4 Performance, robustness and validation evidence compiled and traceable to requirements, risks and intended purpose. T FP
R10.5: Compliance evidence mapping ISO/IEC 42001 (§§  6.1, 8.1––8.4, 9.1–9.3, 10); ISO/IEC 23894 (§§ 5.3–5.8); ISO/IEC 27001; ISO/IEC 27005 M10.5 Structured compliance mapping demonstrating how Annex IV and Article 17 obligations are satisfied. C P
@@ 3.11. R11: Record-keeping
Article 18 of the AI Act mandates that providers retain automatically generated logs and relevant records to ensure traceability and accountability throughout the AI lifecycle. Record-keeping supports post-market monitoring, facilitates incident investigation, and provides evidence for conformity verifications. It must include documentation of model versions, dataset lineage, and data retention or deletion policies, consistent with organisational governance and data protection requirements, including GDPR.
Requirement References Methods Type Target
@@
R11.1: Operational logs ISO/IEC 42001 (§§  7.5, 8.3, 9.1); ISO/IEC 24970 M11.1 Operational logging system in place with records supporting Art. 18 record-keeping and PMM requirements. C FP
@@
R11.2: Version control records ISO/IEC 42001 (§§  7.5, 8.3) M11.2 Version control ensures full traceability and reproducibility of AI system states over time. C D&M
@@
R11.3: Model & dataset lineage ISO/IEC 42001 (§§  7.5, 8.3); ISO/IEC 23894 (§§ 5.3–5.8) M11.3 Full lineage enables auditing, incident analysis, risk traceability and evidence-based updates. C D&M
@@
R11.4: Retention & deletion policies ISO/IEC 42001 (§§  7.5, 8.3); ISO/IEC 23894 (§§ 5.3–5.8); ISO/IEC 27001; ISO/IEC 27005; GDPR (Art. 5.1.c, e) M11.4 Retention/deletion policies implemented to ensure compliance, minimise risk and maintain only necessary records, in line with GDPR data minimisation and storage limitation principles C P
4. Case Study
This section illustrates how the mapping proposed in Section 3 can be applied in practice through a real, ongoing high-risk AI use case currently being analysed and tested in collaboration with Scania, a Swedish automotive manufacturer. While the mapping is generic and technology-agnostic, this example shows how it can structure assurance activities across the AI lifecycle, from design-time verification to validation and documentation. Table LABEL:tab:use_case presents the results of this mapping process; for the sake of brevity, it is reported only partially and serves as an illustrative example. The use case concerns an AI-based system for detecting cyberattacks within connected vehicles in an industrial setting at Scania. Modern vehicles rely on internal communication networks that enable real-time data exchange between electronic components. Protecting these networks is critical, as cyberattacks may compromise both vehicle safety and system reliability. To address these risks, automotive manufacturers increasingly deploy Intrusion Detection Systems (IDS) that monitor in-vehicle communications and identify abnormal or potentially harmful behaviour. In this setting, the IDS employs AI techniques to analyse in-vehicle network traffic and detect patterns associated with known or emerging attacks. The system processes network data, extracts relevant features, and uses machine-learning models to distinguish normal operation from suspicious activity. To support transparency and engineering oversight, explainability mechanisms are integrated to allow engineers to understand why specific activities are flagged as intrusions. Designed for real-time deployment in vehicle control units or central gateways, this use case demonstrates the application of the proposed mapping in a complex, safety-critical automotive context. The mapping enables structured and traceable verification of data handling, model behaviour, explainability, and documentation, providing evidence to support safety, cybersecurity, and regulatory requirements without intervening in system development.
Req. Sub-req. Method Instantiation in the Use Case
@@ 4. Case Study
5. Discussion
Rather than functioning solely as an analytical construct, the proposed mapping is intended to support the institutional ecosystem through which AI Act compliance is operationalised in practice. By linking legal requirements to identifiable and verifiable assessment activities, the mapping provides a common reference structure that can be reused across provider self-assessment, regulatory sandbox experimentation, and formal conformity assessment. This addresses a central implementation risk of the AI Act: that identical legal obligations may be translated into materially different verification practices across organisations, sectors, and Member States.
For providers, the mapping enables structured self-assessment as part of continuous monitoring and post-market obligations. By making explicit which verification activities correspond to which legal requirements, it supports systematic internal evaluations over time, including reassessment following model updates, data drift, or changes in deployment context. This reduces reliance on implicit or ad hoc interpretations of compliance and facilitates the generation of traceable evidence that can be reused across internal governance, regulatory reporting, and external review. In regulatory sandboxes established under Article 57, the mapping can function as a shared interpretive layer between developers and competent authorities. By clarifying how experimental testing activities relate to specific AI Act obligations, it supports the structuring of testing plans, documentation practices, and exit reports, while preserving the flexibility required for innovation-oriented assessment. This role is particularly relevant in a cross-border context, where divergent sandbox practices risk emerging in the absence of a common operational reference. European-funded initiatives such as AI Factories (eu_ai_factories) and Testing and Experimentation Facilities (TEFs) (eu_tefs) are expected to play an increasingly central role in supporting the testing and validation of AI systems, including in the context of regulatory sandboxes, as reflected in recent policy and implementation guidance (implementingAct). In this setting, the availability of a shared mapping between legal requirements and verification activities becomes critical. These infrastructures are not merely technical providers, but intermediaries that translate regulatory expectations into concrete testing capabilities. A common mapping helps ensure that testing performed within such initiatives is interpretable, comparable, and reusable across regulatory and institutional contexts, rather than being tied to local or project-specific interpretations.
More broadly, the mapping contributes to convergence in verification methodologies across the EU by offering a reusable, technology-agnostic operational vocabulary. While enforcement remains decentralised and sectoral specificities must be respected, convergence on underlying verification logic is essential to avoid fragmentation of compliance practices. The mapping is therefore intentionally designed to be extensible: it can be refined, extended, and instantiated by different actors, including competent authorities, notified bodies, EU-funded infrastructures, and sectoral initiatives. Such extensions can introduce domain-specific metrics or procedures while preserving traceability to the original legal requirements. In this sense, the mapping is best understood as an intermediate coordination layer between legal text and assessment practice. It does not prescribe how verification must be performed, but provides a common structure within which diverse practices can evolve in a coherent and interoperable manner. Encouraging its extension and sectoral instantiation is therefore not a limitation, but a necessary condition for achieving both regulatory consistency and practical relevance under the AI Act.
Finally, while this work is grounded in the European regulatory context and explicitly targets the implementation of the AI Act, the underlying challenge it addresses is not unique to Europe. Many jurisdictions are currently developing or refining AI-specific regulatory frameworks that similarly rely on high-level, principle-driven obligations whose practical verification remains underspecified. In this respect, the approach adopted in this paper is intended to be transferable. The notion of an explicit mapping between legal requirements and verifiable assessment activities can be adapted to other regulatory settings, supporting regulatory learning, comparability, and institutional capacity-building beyond the EU. The authors therefore encourage the development of analogous mappings in other jurisdictions and see value in future cross-jurisdictional dialogue on verification methodologies as AI regulation continues to evolve globally.
6. Conclusion
This paper introduced an explicit and structured mapping between high-level obligations under the EU AI Act and concrete, verifiable assessment activities. Starting from 11 high-level requirements used as a normative basis, the mapping decomposes these obligations into 48 operational sub-requirements and associates them with 66 verification activities. By making these relationships explicit, the mapping bridges the gap between regulatory intent and assessable technical and organisational evidence across the AI lifecycle.
A central contribution of this work is not merely the enumeration of verification activities, but the articulation of a reusable mapping logic. By organising verification activities along two orthogonal dimensions, i.e. the type of verification performed and the lifecycle target to which it applies, the mapping makes visible how compliance with the AI Act is constructed in practice. It shows that conformity emerges from structured combinations of procedural controls and empirical testing, rather than from isolated checks or single metrics. By formalising these relationships, the mapping directly reduces key sources of uncertainty in AI Act implementation. Interpretive uncertainty is addressed by clarifying how abstract legal obligations can be decomposed into operational elements grounded in recognised practices. Operational uncertainty is reduced by identifying verification activities that are implementable, repeatable, and auditable across different organisational and technical contexts. Procedural uncertainty is mitigated by enabling consistent documentation, traceability, and comparison of evidence across systems, actors, and regulatory settings.
The value of the mapping extends beyond individual assessments. It provides a common reference that can be reused across provider self-assessment, regulatory sandboxes, conformity assessment, and post-market monitoring, supporting convergence in verification practices across Member States while preserving contextual flexibility. In doing so, it aligns with the AI Act’s governance model, which relies on continuous oversight and regulatory learning rather than one-off certification. The mapping is intentionally technology-agnostic and non-prescriptive. Its purpose is not to fix verification practices in advance, but to offer a stable coordination layer between legal text and assessment practice that can be extended, refined, and instantiated by different actors, including sectoral initiatives and EU-funded testing infrastructures. In this sense, the contribution of this work lies in establishing a shared operational grammar for AI Act compliance verification, one that enables consistency without rigidity and innovation without fragmentation.
Acknowledgement
This work is supported by the EU project Citcom.AI and Vinnova INTERSTICE project (reference number: 2024-00661). This work is also supported by the Swedish AI Factory and the Luxembourg AI Factory.